OpenStudy (anonymous):

Write a recursive formula for the sequence 15, 26, 48, 92, 180, …. Then find the next term.

OpenStudy (anonymous):

Which series is it?

OpenStudy (anonymous):

well, i thnk its gunna start with an=

OpenStudy (anonymous):

\[a _{n+1} = 2a _{n} - 4, \forall n \ge1 \]

OpenStudy (anonymous):

See the difference between every two numbers you will get diff as 11,22,44 etc

OpenStudy (anonymous):

A. an = 2an – 1 – 4 where a1 = 15; 356 B. an = 2an – 4 where a1 = 15; 356 C. an = 4 + 11 • 2n – 1 where a1 = 15; 356

OpenStudy (anonymous):

should be A